Zari:

The leg of a right triangle is 3 units and the hypotenuse is 7 units. What is the length, in units, of the other leg of the triangle? 4 Square root of 40 Square root of 58 10

axie:

Ok so what do you know about the triangle side length formula?

Zari:

i think its 10 or 4

axie:

Leg2+Leg2=Hypotenuse2 In this case we have 3^2 + Leg2 = 7^2

Zari:

yeah

axie:

So how would you solve this?

Zari:

7+2 ?

axie:

No not quite

Zari:

oh

axie:

Ok first we do 3^2, this is 9 bc its 3x3. Second we do 7^2, this is 49 or 7x7 So we have this 9+Leg^2=49

axie:

what would you do from here?

Zari:

@axie wrote:
Ok first we do 3^2, this is 9 bc its 3x3. Second we do 7^2, this is 49 or 7x7 So we have this 9+Leg^2=49 ohhhh ok

Zari:

is it 40?

axie:

To simplify, we would do 49-9=40 we do this so that we can get rid of 9 on one side of the equation so we have leg^2=40 what from here?

Zari:

okay

axie:

From there we take the square root of 40 but that's a desmol making your answer the square root of 40

Zari:

ok